OpenStudy (anonymous):

Fabian has a savings account balance of $1,374.73. The interest rate on the account is 3.9% compounded twice a year. If he opened the account nine years ago, what was the value of his initial deposit?

OpenStudy (anonymous):

\[A = P(1 + \frac{r}{n})^{nt}\]\[1374.73 = P(1 + \frac{0.039}{2})^{2(9)}\]Solve for P which is the initial amount.